Kellen​'s boat travels 12 mph. Find the rate of the river current if she can travel 6mi upstream in the same amount of time she can go 10mi downstream.(let x = rate of the current)
Answer by addingup(3677)

distance = rate*time
Let x = rate of the current (this is what the problem says)
6/(12-x) = 10/(12+x)
cross multiply:
72+6x = 120-10x
16x = 48
x = 3 this is the speed of the current.